# Dual-Level Boost Network for Long-Tail Prohibited Items Detection in X-ray Security Inspection

### v1

@@ 3.4 Model-level Related Proposals Aggregation
For each of the top-k proposals, let the feature vector extracted from each proposal be denoted as $f_{i_{j}}$, where $f_{i_{j}}$ is the feature vector associated with proposal $p_{i_{j}}$, $j=1,2,\dots,k$. We concatenate these feature vectors to form a single concatenated feature representation:
@@
where the feature $F_{\text{concat}}$ captures the relationships and dependencies among the top-$k$ proposals, encoding contextual information that reflects prohibited item interactions. To process the concatenated feature vector $F_{\text{concat}}$, we apply a fully connected layer to create a fused relational feature representation. Let $\mathbf{W}_{1}$ and $\mathbf{b}_{1}$ represent the weights and bias of the fully connected layer, respectively. The relational fusion feature $F_{\text{fusion}}$ can be computed as follows:
$F_{\text{fusion}}=\sigma(\mathbf{W}_{1}F_{\text{concat}}+\mathbf{b}_{1})$ (9)
where $\sigma$ represents an activation function, such as ReLU. This fused feature vector $F_{\text{fusion}}$ encodes higher-order relationships among the top-k proposals, integrating contextual information across all the selected prominent objects. The fused relational feature $F_{\text{fusion}}$ is then concatenated with each of the original top-k proposal features to further enhance their individual representations. For each proposal feature $f_{i_{j}}$, we form an augmented feature vector $f_{i_{j}}^{\text{aug}}$ by concatenating with $F_{\text{fusion}}$:
$f_{i_{j}}^{\text{aggr}}=\text{Concat}(f_{i_{j}},F_{\text{fusion}})$ (10)
To match the dimensionality of the original proposal features, a second fully connected layer with weights $\mathbf{W}_{2}$ and bias $\mathbf{b}_{2}$ is applied to each aggregated feature $f_{i_{j}}^{\text{aggr}}$, producing a new feature $f_{i_{j}}^{\text{new}}$ for each proposal:
$f_{i_{j}}^{\text{new}}=\sigma(\mathbf{W}_{2}f_{i_{j}}^{\text{aggr}}+\mathbf{b}_{2})$ (11)
This operation ensures that $f_{i_{j}}^{\text{new}}$ has the same dimensions as the original proposal feature $f_{i_{j}}$, allowing it to seamlessly replace the original in subsequent processing steps. Worth mentioning that the use of an activation function depends on whether the specific detection heads apply an activation function during feature extraction and score generation, as we need to align the distribution of the features and predicted values.
The final step involves substituting the original classification logits—computed from the initial proposal features—with the updated features $f_{i_{j}}^{\text{new}}$. These updated features, now enriched with relational context from other high-confidence proposals, serve as the inputs for the classification and bounding box regression heads. Let $\text{Logits}_{\text{new}}$ represent the classification logits computed from the new feature representations:
@@
This replacement completes the augmentation process, equipping the model with enhanced, context-aware feature representations that improve detection performance, particularly these tail classes in cluttered or occluded scenes.
@@ 3.5 Overall Training Process
Algorithm 1 summarizes the whole process. This algorithm presents the training procedure for DBNet, summarizing the key stages of data-level and model-level modules, and training. The data-level module involves instance cropping and Poisson fusion to generate synthetic data for tail classes. The model-level module enhances the feature representations of proposals using relational context derived from high-confidence proposals. Finally, the detection model is trained using the augmented dataset and context-enriched features. DBNet, through the integration of data-level and model-level modules, achieves superior performance, particularly for tail classes, by addressing the challenges of long-tail distributions in X-ray prohibited items detection.
@@
Output: Trained model $\mathbf{M}$, augmented dataset $D^{\prime}$.
Data-Level Augmentation:
@@
Model-Level Aggregation:
@@
Training:
Train detection model $\mathbf{M}$ using augmented dataset $D^{\prime}$ and updated features.
Algorithm 1 Training Procedure of DBNet
@@ 4 Experiments
In this section, we present a comprehensive evaluation of the proposed DBNet. Section 4.1 introduces the experimental setup. In Section 4.2, we assess our model’s effectiveness and generalizability by integrating it into various baseline detection models. Section 4.3 compares our method against sota long-tail detection techniques. Finally, in Section 4.4, we conduct ablation studies to validate the contribution of each data-level and model-level module and analyze the impact of the parameter $k$ of the model-level.
Method HiXray Dataset PIDXray Dataset
@@
Table 1: Comparison of detection performance between baseline models and DBNet-enhanced (+ours) versions on the HiXray and PIDXray datasets. The proposed DBNet consistently improves AP50 of the overall categories, especially for these tail ones in HiXray dataset.
@@ 4.1 Experimental Setup
Tasks and Datasets. To ensure fair comparisons with previous detection methods, we evaluate our model on two large-scale X-ray datasets, HiXray [34] and PIDXray [39]. HiXray is derived from real-world scenarios and features a distinctly long-tail distribution, making it ideal for evaluating model performance on naturally imbalanced data. Although PIDXray samples are synthetically generated and balanced, this dataset offers a comprehensive evaluation framework for X-ray imaging tasks and is suitable for testing the classification capabilities of the proposed DBNet.
@@
Implementation Details. All the experiments were conducted using the mmdetection open-source toolbox [^2] https://github.com/open-mmlab/mmdetection [5] to ensure consistency and comparability. The parameter settings for all models followed the default configurations of mmdetection. Experiments were conducted on two NVIDIA GeForce RTX 4090 GPUs.
@@ 4.2 Comparison with Various Detection Baselines
To determine the performance and generalizability of DBNet, we integrated it with multiple popular object detection frameworks, including F-RCNN (Faster RCNN [26]), S-RCNN (Sparse RCNN [31]), RetinaNet [27], CenterNet [9], and ATSS [45]. The results are summarized in Table 1.
@@
Performance on HiXray Dataset. Table 1 shows that DBNet consistently improved detection accuracy across all baseline models on the HiXray dataset. For Faster RCNN, the baseline achieved an $\text{AP}_{50}$ of 84.2%, whereas integrating DBNet increased it to 85.2%. The most significant gains were observed in tail classes like class NL, which saw an improvement from 32.0% to 41.7%, an increase of 9.7%. This demonstrates DBNet’s strength in boosting the performance of tail classes. For ATSS, DBNet provided an $\text{AP}_{50}$ of 85.6%, compared to the baseline’s 83.5%. Specific improvements included: (1) Class NL: Improved from 28.3% to 40.7%, which is crucial given that NL is a challenging tail class with limited representation. (2) Class PO2: AP increased from 93.6% to 94.2%, which, although modest, shows improved consistency in handling frequent classes. Overall, the AP50 for ATSS improved by 2.1%, demonstrating that DBNet’s dual-level augmentations effectively balance the performance between head and tail classes.
@@
Performance on PIDXray Dataset. The DBNet also demonstrated strong performance gains on the PIDXray dataset. Faster RCNN, for instance, saw an increase in AP50 from 76.0% to 77.6%. Notable improvements include: (1) The AP of class GU improved from 47.6% to 48.2%. This improvement, while relatively small, demonstrates DBNet’s ability to provide more robust features for objects that may be occluded or embedded within a complex context. Guns, in particular, often have parts that overlap with other objects, and the improvement suggests better discrimination for such features. (2) The AP of class SP rose from 67.3% to 72.8%. This significant increase indicates that DBNet’s augmentations made the model better at recognizing screwdrivers even when partially occluded or in cluttered environments, which is a common scenario in X-ray imagery. Across all baseline models, DBNet’s integration led to increased detection performance for tail classes, highlighting the robustness of both data-level and model-level augmentation strategies. For instance, using S-RCNN with DBNet on the PIDXray dataset resulted in an increase in AP for GU from 37.1% to 38.5%, which is particularly notable given the scarcity of instances for this class. This shows DBNet’s capability in enhancing detection for underrepresented items by increasing presence and importance.
@@
Table 2: Comparison with popular long-tail detection methods on the HiXray dataset. Results highlight the advantage of DBNet in addressing data imbalance, particularly for rare prohibited items.
4.3 Comparison with Long-Tail Detection Methods
We compared DBNet with several sota long-tail task methods including Seesaw Loss [40], GHM Loss [17], Focal Loss [27], and standard Cross Entropy (CE) [46]. The results of performance comparison is presented in Table 3.
Comparison Analysis on HiXray Dataset. DBNet outperformed all other methods in terms of AP50, specifically focusing on tail class improvement. The AP50 for Seesaw Loss was 83.8%, while DBNet achieved 85.0%, a significant improvement of 1.2%. Some specific examples of AP improvements include: (1) The AP of class NL increased from 35.5% using Seesaw Loss to 41.2% with DBNet. This improvement demonstrates that the data-level augmentation component in DBNet effectively balanced the representation of tail classes, which Seesaw Loss alone was unable to achieve at the same level. (2) The AP50 of CO reduced from 69.5% using Seesaw Loss to 67.4% using DBNet. Although there was a slight reduction in AP50 for this class, DBNet maintained competitive performance, indicating that its augmentations did not lead to significant regressions for head classes while still boosting the overall tail class.
Consistent Gains Across Long-Tail Methods. When compared with Focal Loss, DBNet consistently provided higher AP values for classes with fewer samples. For example: (1) The AP50 of class TA improved from 91.5% to 94.3%. This improvement suggests that the model-level augmentation in DBNet, which enhances the contextual relationships among detected objects, was particularly effective in aiding the detection of tape, which is often surrounded by other objects in cluttered settings. (2) The AP50 of Class PO1 increased from 94.7% with Focal Loss to 95.3% using DBNet. The gain shows that DBNet is better at handling challenging objects by making more informed decisions based on both object-level features and their spatial relationships. These improvements demonstrate that the proposed DBNet effectively addresses the long-tail distribution, particularly with severe class imbalances.
@@ 4.4.1 Effectiveness of Each Module
The ablation results of each module are provided in Table 3. We analyzed the effects of enabling data-level augmentation, model-level augmentation, and both.
@@ D M PO1 PO2 WA LA MP TA CO NL
Table 3: Ablation study on individual contributions of data-level and model-level modules. The table demonstrates how each component contributes to the overall detection accuracy improvement.
As illustrated in Table 3, Baseline with no augmentation achieved an AP50 of 84.2%. This served as our control, providing a point of comparison to assess the individual and combined contributions of each module. Data-Level Augmentation only improved to 85.0%. The class-specific performance showed notable gains for tail classes, such as Class NL which increased from 32.0% to 41.2%. This increase illustrates how data-level augmentation directly addresses data scarcity issues by providing more diverse training samples for underrepresented classes. Model-Level Strategy only resulted in an AP50 of 84.6%. The smaller overall improvement compared to data-level augmentation suggests that model-level augmentation is more complementary rather than a complete solution by itself. Combined Augmentation with full DBNet achieved the best performance, with an AP50 of 85.2%. Notably, Class PO1 improved from 95.0% to 96.1%, and Class PO2 increased from 93.4% to 94.4%. These improvements confirm that both augmentation strategies are highly effective when combined, each addressing different aspects of the long-tail distribution problem.

### v2

@@ 3.3 Contextual Feature Integration
For each of these top-$k$ proposals, we extract their feature vectors, denoted as $f_{i_{j}}$ for $j=1,2,\dots,k$. To capture the relationships among these top proposals, we concatenate their feature vectors into a single vector:
@@
This concatenated feature vector $F_{\text{concat}}$ contains information about the dependencies between the top-$k$ proposals, reflecting how they interact in the context of the image. We then apply a fully connected layer to process $F_{\text{concat}}$ and produce a relational fusion feature, $F_{\text{fusion}}$, as follows:
$F_{\text{fusion}}=\sigma(\mathbf{W}_{1}F_{\text{concat}}+\mathbf{b}_{1}),$ (4)
where $\sigma$ is an activation function, such as ReLU, and $\mathbf{W}_{1}$ and $\mathbf{b}_{1}$ are the weights and bias of the fully connected layer. The relational feature $F_{\text{fusion}}$ captures the higher-order interactions between the selected proposals, encoding the contextual relationships that are crucial for distinguishing between visually similar categories.
To further integrate this relational information, we concatenate the fused relational feature $F_{\text{fusion}}$ with each of the original proposal features $f_{i_{j}}$, forming an augmented feature vector for each proposal. We then apply another fully connected layer to each augmented feature vector to produce the final updated feature vector $f_{i_{j}}^{\text{new}}$. This process can be formulated as follows:
$f_{i_{j}}^{\text{aug}}=\text{Concat}(f_{i_{j}},F_{\text{fusion}}),$ (5)
$f_{i_{j}}^{\text{new}}=\sigma(\mathbf{W}_{2}f_{i_{j}}^{\text{aug}}+\mathbf{b}_{2}),$ (6)
where $\mathbf{W}_{2}$ and $\mathbf{b}_{2}$ are the weights and bias of the second fully connected layer. This updated feature vector $f_{i_{j}}^{\text{new}}$ incorporates both the original features of the proposals and the relational information from other proposals, enabling the model to make more informed classification decisions.
Finally, the updated features $f_{i_{j}}^{\text{new}}$ replace the original proposal features in the classification and bounding box regression heads. The new classification logits are computed from these enhanced features:
@@ 3.4 Overall Training Process
Algorithm 1 outlines the entire training procedure for XIDNet, highlighting the key stages of the X-ray-Specific augmentation and Contextual Feature Integration, along with the overall training process. The X-ray-Specific augmentation module involves cropping instances and applying Poisson fusion to generate synthetic data for tail classes. The Contextual Feature Integration module improves feature representations of proposals by leveraging relational context from high-confidence proposals. Finally, the detection model is trained using the augmented dataset and context-enhanced features. By integrating these two modules, XIDNet addresses the long-tail distribution problem in X-ray prohibited items detection, delivering superior performance, especially for tail categories.
@@
Output: Trained model $\mathbf{M}$.
X-ray-Specific Augmentation:
@@
Contextual Feature Integration:
Sample Batches from the $D^{\prime}$;
@@
Perform backpropagation and update parameters;
Output the trained model $\mathbf{M}$.
Algorithm 1 Training Procedure of XIDNet
@@ 4 Experiments
We provide a comprehensive evaluation of XIDNet. Sec. 4.1 outlines the experimental setup. Sec. 4.2 evaluates the effectiveness and generalizability of integrating it into various baseline models. Sec. 4.3 compares XIDNet with SoTA long-tail detection techniques, while Sec. 4.5 contrasts it with popular X-ray detection methods. Sec. 4.6 presents ablation studies to assess the contributions of each module, as well as the impact of the parameter $k$. Finally, Sec. 4.7 includes experiments on manual segmentation scalability and generalization to transformer-based models.
Method HiXray Dataset PIDray Dataset
@@
Table 1: Comparison of detection performance between baselines and XIDNet-enhanced (+ours) versions on the HiXray and PIDray datasets. The proposed XIDNet consistently improves AP50 of the overall categories, especially for these tail ones in HiXray dataset.
@@ 4.1 Experimental Setup
Tasks and Datasets. To ensure fair comparisons with previous detection methods, we evaluate our model on two large-scale X-ray datasets, HiXray [32] and PIDray [36]. HiXray is derived from real-world scenarios and features a distinctly long-tail distribution, making it ideal for evaluating model performance on naturally imbalanced data. Although PIDray samples are synthetically generated and balanced, this dataset offers a comprehensive evaluation framework for X-ray imaging tasks and is suitable for testing the classification capabilities of the proposed XIDNet.
@@
Implementation Details. All the experiments were conducted using the mmdetection open-source toolbox [^1] https://github.com/open-mmlab/mmdetection [5] to ensure consistency and comparability. The parameter settings for all models followed the default configurations of mmdetection. Experiments were conducted on four NVIDIA GeForce RTX 4090 GPUs.
@@ 4.2 Comparison with Various Detection Baselines
To determine the performance and generalizability of XIDNet, we integrated it with multiple popular object detection frameworks, including F-RCNN (Faster RCNN [25]), S-RCNN (Sparse RCNN [29]), RetinaNet [26], CenterNet [8], and ATSS [44]. The results are summarized in Tab. 1.
@@
Performance on HiXray Dataset. Tab. 1 shows that XIDNet consistently improved detection accuracy across all baseline models on the HiXray dataset. For Faster RCNN, the baseline achieved an $\text{AP}_{50}$ of 84.2%, whereas integrating XIDNet increased it to 85.2%. The most significant gains were observed in tail classes like class NL, which saw an improvement from 32.0% to 41.7%, an increase of 9.7%. This demonstrates XIDNet’s strength in boosting the performance of tail classes. For ATSS, XIDNet provided an $\text{AP}_{50}$ of 85.6%, compared to the baseline’s 83.5%. Specific improvements included: (1) Class NL: Improved from 28.3% to 40.7%, which is crucial given that NL is a challenging tail class with limited representation. (2) Class PO2: AP increased from 93.6% to 94.2%, which, although modest, shows improved consistency in handling frequent classes. Overall, the AP50 for ATSS improved by 2.1%, demonstrating that XIDNet’s dual-level augmentations effectively balance the performance between head and tail classes.
@@
Performance on PIDray Dataset. The XIDNet also demonstrated strong performance gains on the PIDray dataset. Faster RCNN, for instance, saw an increase in AP50 from 76.0% to 77.6%. Notable improvements include: (1) The AP of class GU improved from 47.6% to 48.2%. This improvement, while relatively small, demonstrates XIDNet’s ability to provide more robust features for objects that may be occluded or embedded within a complex context. Guns, in particular, often have parts that overlap with other objects, and the improvement suggests better discrimination for such features. (2) The AP of class SP rose from 67.3% to 72.8%. This significant increase indicates that XIDNet’s augmentations made the model better at recognizing screwdrivers even when partially occluded or in cluttered environments, which is a common scenario in X-ray imagery. Across all baseline models, XIDNet’s integration led to increased detection performance for tail classes, highlighting the robustness of both data-level and model-level augmentation strategies. For instance, using S-RCNN with XIDNet on the PIDray dataset resulted in an increase in AP for GU from 37.1% to 38.5%, which is particularly notable given the scarcity of instances for this class. This shows XIDNet’s capability in enhancing detection for categories with insufficient training samples.
4.3 Comparison with Long-Tail Detection Methods
We compared XIDNet with several sota long-tail task methods. The methods considered in this comparison are LogN [47], ROG [45], Seasaw [37], GHM [19], and Focal [26]. These methods are popular for addressing the challenges posed by class imbalances, especially for the tail category.
XIDNet outperforms all the methods in terms of overall detection performance, achieving an AP50 score of 85.2, surpassing ROG (84.7) and Seasaw (83.8). The proposed framework integrates both X-ray-specific augmentation and contextual feature integration, making it particularly effective in overcoming the challenges posed by long-tail distributions. These innovations allow XIDNet to not only perform well on head categories but also provide significant improvements for tail categories, where most other methods are struggle to achieve the balance.
@@ 4.3 Comparison with Long-Tail Detection Methods
Table 2: Comparison with popular long-tail detection methods on the HiXray dataset. Results highlight the advantage of XIDNet in addressing data imbalance, particularly for rare prohibited items.
Focusing on head categories, XIDNet maintains strong performance, achieving 96.1 in PO1 and 94.4 in PO2. This is higher than ROG, which scores 95.8 and 94.0, and Seasaw, which achieves 95.4 and 94.3. This indicates that XIDNet is on par with state-of-the-art methods in handling the majority classes. However, its true strength lies in the detection of tail categories, particularly in the detection of NL category, where XIDNet scores 41.7, a substantial improvement over ROG (35.2) and Seasaw (35.5). This highlights XIDNet’s ability to address the severe class imbalance typically encountered in X-ray prohibited item detection.
The performance of XIDNet in tail category detection is significantly better than conventional methods, particularly GHM (30.1) and Focal (30.2), which have a much lower AP50 in the tail class. XIDNet’s ability to generate synthetic data for tail classes through Poisson blending and its integration of contextual information from surrounding objects in the image contribute to its superior performance. This combined approach leads to enhanced feature representations for tail items, allowing the model to better handle occlusions and complex scenes. Therefore, XIDNet not only excels in detecting common prohibited items but also proves to be highly effective for rare items, making it a robust solution for X-ray screening in security applications.
4.4 Comparison with Data Augmentation Methods
In this section, we compare the performance of XIDNet with several popular data augmentation methods, including Copy-paste [11], Mixup [42], and Cutout [7]. These methods are widely used for enhancing the performance of object detection models, especially in the context of imbalanced datasets. The results are illustrated in Tab. 3.
The Copy-paste augmentation method achieves an impressive average AP50 of 84.4, with particularly strong performance across the head categories (PO1, PO2) and moderate performance for tail categories. However, as expected, its performance drops significantly in tail categories such as NL (26.4), highlighting the challenge of handling rare classes with this method. For fairness, we also used manually optimized instances in the Copy-paste experiment. Mixup yields an AP50 of 80.8, which is lower than Copy-paste. It performs well in head categories, particularly PO1 (94.4) and PO2 (92.8), but similarly struggles with tail categories, with a relatively low performance for NL (16.3). Cutout provides an AP50 of 83.9, ranking between Copy-paste and Mixup in overall performance. Its performance in head categories is competitive (e.g., PO1: 95.5, PO2: 94.0), but, like other augmentation methods, it experiences a significant drop in performance for tail category (NL: 25.2).
All data augmentation methods perform well in head classes (PO1, PO2) and achieve high AP50 scores for these categories. Copy-paste leads with the highest overall AP50 and performs particularly well across multiple head categories. The most significant challenge for these augmentation methods is improving the performance of tail categories (e.g., NL). Copy-paste achieves the best performance in this category compared to Mixup and Cutout, but it still falls short of what is needed for a robust tail-class detection.
Method AP50 AP50 across various categories
PO1 PO2 WA LA MP TA CO NL
Copy-paste [11] 84.4 96.3 95.7 94.1 98.0 98.2 96.4 69.7 26.4
Mixup [42] 80.8 94.4 92.8 89.3 97.9 96.9 94.2 64.8 16.3
Cutout [7] 83.9 95.5 93.9 91.6 98.2 97.5 96.2 73.2 25.2
DOAM[39] 82.2 95.9 94.7 93.7 98.1 98.1 95.8 65.0 16.1
LIM [32] 83.2 96.1 95.1 93.9 98.2 98.3 96.4 65.8 21.3
XIDNet (ours) 83.5 94.5 93.3 92.0 98.1 97.2 94.5 67.7 31.0
Table 3: Comparison with popular augmentation methods and X-ray-specific detection methods on HiXray (base model: ATSS).
4.5 Comparison with X-ray Detection Methods
In this section, we compare the performance of XIDNet with two popular X-ray prohibited items detection methods: DOAM [39] and LIM [32], which are specifically designed for X-ray images. The results are illustrated in Tab. 3.
The DOAM method achieves an AP50 of 82.2, showing strong performance for head categories (PO1: 95.9, PO2: 94.7). However, it suffers from significant degradation in performance for tail categories, especially NL (16.1), highlighting the difficulty of handling the categories with insufficient training samples. Moverever, LIM outperforms DOAM, with an AP50 of 83.2. It performs similarly in head categories, achieving strong results across PO1 (96.1), PO2 (95.1), and others, but still faces challenges with tail categories, such as NL (21.3), though its performance in tail classes is slightly better than the DOAM method.
Both DOAM and LIM exhibit strong performance in detecting head categories, with LIM achieving slightly better results overall. However, neither method achieves the robustness needed to handle tail classes effectively. XIDNet outperforms both DOAM and LIM in tail class performance, with a significantly higher AP50 of 31.0 in the NL category. This demonstrates the effectiveness of XIDNet’s data augmentation strategies and the contextual feature integration module, which are designed to improve detection accuracy in cluttered or occluded scenes, especially for the categories with insufficient training samples.
@@ 4.6.1 Effectiveness of Each Module
The ablation results of each module are provided in Tab. 5. We analyzed the effects of enabling data-level augmentation, model-level augmentation, and both.
@@ D M PO1 PO2 WA LA MP TA CO NL
Table 4: Ablation study on individual contributions of data-level (D) and model-level (M) modules. The table demonstrates how each component contributes to the overall detection improvement.
As illustrated in Tab. 5, Baseline with no augmentation achieved an AP50 of 84.2%. This served as our control, providing a point of comparison to assess the individual and combined contributions of each module. Data-Level Augmentation only improved to 85.0%. The class-specific performance showed notable gains for tail classes, such as Class NL which increased from 32.0% to 41.2%. This increase illustrates how data-level augmentation directly addresses data scarcity issues by providing more diverse training samples for underrepresented classes. Model-Level Strategy only resulted in an AP50 of 84.6%. The smaller overall improvement compared to data-level augmentation suggests that model-level augmentation is more complementary rather than a complete solution by itself. Combined Augmentation with full XIDNet achieved the best performance, with an AP50 of 85.2%. Notably, Class PO1 improved from 95.0% to 96.1%, and Class PO2 increased from 93.4% to 94.4%. These improvements confirm that both augmentation strategies are highly effective when combined, each addressing different aspects of the long-tail distribution problem.